in the incoming year, the probability is 0.30 that the price of a certain house will go up 20%, the probability is 0.50 that the price will go up 10%, the probability is 0.10 that the price will remain the same, and the probability is 0.10 that the price will go down 10%. find the expected percentage increase.
answer please.....

Answer by Theo(13342)

.3 * 20% + .5 * 10% + .1 * 0% + .1 * -10% = 10%

the expected percentage increase is 10%.

you can also calculate the expected percent increase as follows:

assume the value of the investment is x.

30% of the time it will increase 20% to become 1.2 * x
50% of the time it will increase 10% to become 1.1 * x
10% of the time it will increase 0% to become 1.0 * x
10% of the time it will decrease 10% to become .9 * x

expected value of the investment becomes:

.3 * (1.2 * x) + .5 * (1.1 * x) + .1 * (1.0 * x) + .1 * (.9 * x)

you will get expected value = 1.1 * x.

the percent increase in value is 1.1 * x / 1.0 * x = .1.

multiply that by 100 and it becomes 10%.